# On the Federated Learning Framework for Cooperative Perception

### v1

@@ IV Experiments and Analysis
The experiments were conducted on a computer equipped with a single NVIDIA GeForce RTX 4090 GPU and a 12th Generation Intel Core i9-12900K CPU. The software environment consisted of Pytorch 2.0.1 and CUDA 12.3.
@@ IV-A Experiment Setup
Our experimental evaluation leveraged the extensive and publicly available CP datasets from OPENV2V [29], which were augmented with an additional dataset as detailed in [27]. This enhancement broadens the diversity and volume of the scenarios represented, providing a robust foundation for testing our model.
@@
Following the methodology established in [27], our study orchestrates a collaborative framework among four distinct industrial clients to assess the efficacy of our proposed object detection model under various operational conditions. Specifically, the dataset comprises:
@@
These entities collaboratively trained their models using a federated learning approach, which improved their perception capabilities. This arrangement ensured that while each entity benefited from shared data to enhance their models, direct access to the raw sensor data of others was strictly prohibited, addressing privacy concerns.
For a fair comparison between our federated learning strategy and state-of-the-art approaches, our model architecture adheres to the configuration described in [27]. This involves processing camera images observed by clients. Similar to [27], we begin by sending these images through a 3-layer ResNet34 encoder, which is proficient at extracting and encoding image features across various spatial dimensions. This encoding process yields feature tensors at differentiated spatial resolutions: $64\times 64\times 128$, $32\times 32\times 256$, and $16\times 16\times 512$.
@@
Figure 1: IoU Results

### v2

@@ IV Experiments and Analysis
The experiments were conducted on a computer equipped with a single NVIDIA GeForce RTX 4090 GPU and a 12th Generation Intel Core i9-12900K CPU. The software environment consisted of Pytorch 2.0.1 and CUDA 12.3. Our experiments are implemented on BEV semantic segmentation, a crucial task in autonomous driving technologies. BEV semantic segmentation facilitates a comprehensive top-down environmental perspective, which is essential for navigation, obstacle avoidance, and path planning in autonomous vehicles.
@@ IV-A Experiment Setup
We do the experiment on the dataset provided by FedBeVT [27] which incorporates data collected from three distinct types of vehicles: a compact car, a pickup truck, and a bus. These vehicles represent different camera sensor installation positions, which supplements additional data to the OpenV2V dataset [29] to make the dataset more complex. This enhancement broadens the diversity and volume of the scenarios represented, providing a robust foundation for testing our model.
@@
Following the protocol established in [27], our study orchestrates a collaborative framework among four distinct industrial clients to assess the efficacy of our proposed object detection model under various operational conditions Specifically, the dataset comprises:
@@
For a fair comparison between our federated learning strategy and state-of-the-art approach FedBEVT [27], our model architecture adheres to the configuration described in [27]. This involves processing camera images observed by clients. Similar to [27], we begin by sending these images through a 3-layer ResNet34 encoder, which is proficient at extracting and encoding image features across various spatial dimensions. This encoding process yields feature tensors at differentiated spatial resolutions: $64\times 64\times 128$, $32\times 32\times 256$, and $16\times 16\times 512$.
@@
Figure 1: This figure illustrates the results for four distinct clients. The x-axis represents the number of communication rounds, while the y-axis indicates the IoU values. Subfigure (a) displays the performance for the bus client, subfigure (b) shows the results for the truck client, and subfigures (c) and (d) present the outcomes for car client A and car client B, respectively.
@@ IV-B Performance Analysis
The pixel-wise IoU metric is a common evaluation metric for BEV semantic segmentation. It quantifies the overlap between the predicted segmentation and the ground truth, providing a measure of accuracy at the pixel level. The IoU is calculated as follows:
$IoU=\frac{TP}{TP+FP+FN},$ (12)
where TP is the number of true positive pixels, FP is the number of false positive pixels, and FN is the number of false negative pixels.
@@ IV-C Ablation Experiment
Figure 2: This figure is the visualization of model output across different clients. The ego vehicle is on the center. Each panel in this figure presents comparative results for different vehicle clients: (a) Bus, (b) Truck, and (c) Car. For each subfigure, data is organized into six columns. The first column displays camera imagery from the frontal perspective, while the second column shows the rear perspective. The third column represents the ground truth. The fourth column illustrates model output from FedBEVT [27]. The fifth column shows the result from FedDWA without DALoss. The final column depicts result from the model trained with both FedDWA and DALoss, showcasing the effectiveness of the proposed method.